OpenStudy (anonymous):

How would I express the integral ʃsin2x as a power series bounds are from -1 to 1 ?

OpenStudy (unklerhaukus):

\[=\int \sin(2x)\text dx\]\[=-\frac{cos(2x)}{2}+C\]

OpenStudy (unklerhaukus):

\[\cos(z)=\frac{e^{-iz}+e^{iz}}{2}\]

OpenStudy (unklerhaukus):

\[e^x=\sum\limits_{n=0}^∞\frac{ x^n}{n!}\]